OpenStudy (anonymous):

What place in this world can have their temperatures Fahrenheit and Celsius equal?

OpenStudy (anonymous):

it may be some where but i swear by god it is not here !

OpenStudy (anonymous):

siberia is one - northern canada and alasaka probsbly do as well

OpenStudy (anonymous):

-40 C = - 40 F

OpenStudy (anonymous):

yeah buddy i know

OpenStudy (anonymous):

F = 1.8C + 32 x = 1.8x + 32 x - 1.8x = 1.8x + 32 - 1.8x -0.8x = 32 -8x = 320 -x = 40 x = -40 The only place that might be -40 degrees cold must be Antartica.

OpenStudy (anonymous):

no siberia definitely experiences such tempertures

OpenStudy (anonymous):

i like sheg's ans lol

OpenStudy (anonymous):

yea